# Game-Theoretic Regularized Self-Play Alignment of Large Language Models

### v1

@@ 5.1 Effectiveness of Regularization
Table 2: Left: Comparisons of iterative methods with reference models Mistral-7B (Instruct-v0.2). SPPO $\leq$ 9 represents the best results among the $9$ iterations of SPPO. Here, the Regularized SPPO (RSPO) is regularized by the linear combination of Forward KL and Reverse KL divergence, i.e. RSPO (For. + Rev.), where the regularization temperatures are $0.1$ and $0.5$, respectively. Right: Response diversity of SPPO with different regularization methods using Self-BLEU score. The regularization temperatures are the same as in Figure 2 (Right). A lower Self-BLEU score means a higher diversity of the sampled responses. Regularization methods involving Reverse KL resulted in higher diversity of the responses.
Figure 3: Ablation study of regularization temperature $\lambda$ conducted on AlpacaEval 2.0. We evaluate how the average response length and raw WR are affected by the regularization temperature.
5.2 Impact of Different Regularizations
We then study the effect of applying different regularization $R$ in RSPO. To obtain a well-regularized self-play, the tuning of regularization temperature $\lambda$ is necessary. An ablation study of the regularization temperature of different methods is shown in Figure 3. According to the figure, the response length increases along with the temperature in reverse KL divergence and Chi-square divergence regularized RSPO. Meanwhile, the length is decreased with stronger regularization via Forward KL divergence, which was implemented using importance sampling. This result underscores the distinct effects of different regularization strategies. In particular, the raw win rate analysis highlights reverse KL divergence as a crucial factor in enhancing self-play performance. Given that forward KL divergence tends to reduce response length while reverse KL divergence yields significant improvements, we adopt a linear combination of both. This approach is designed to balance their complementary effects, ultimately optimizing for a higher LCWR (RSPO (IS-For. + Rev.) in Figure 2 RHS).
In Figure 2 (Right), we show the results of win rate and LCWR in AlpacaEval 2.0 of different regularizations. Only vanilla Forward KL decreases the win rate of SPPO. The regularizations that consists of Reverse KL including RSPO (Rev. KL) and RSPO (For.+Rev.) have shown significant improvement in win rates compared to vanilla SPPO. In particular, the results of RSPO (For.+Rev.) demonstrate the largest improvement between iterations, achieving the best LCWR.
We study the effect of applying different regularization $R$ in RSPO. In Figure 2 (Right), we show the results of win rate and average response length on AlpacaEval 2.0. Among different regularizations, only Forward KL decreases the win rate of SPPO. The regularizations that consist of Reverse KL including RSPO (Rev. KL) and RSPO (For.+Rev.) have shown significant improvement in win rates compared to vanilla SPPO. In particular, the results of RSPO (For.+Rev.) demonstrate the largest improvement between iterations. We test the best RSPO model on different benchmarks in Table 1, and observe that the proposed method RSPO with proper external regularization achieves superior performance than unregularized self-play and iterative methods like SPPO and iterative DPO, as well as the strong offline alignment method SimPO.  [^2] We report our replicated testing of SPPO Iter3 (https://huggingface.co/UCLA-AGI/Mistral7B-PairRM-SPPO-Iter3) on Arena-Hard. So, it is different from the result presented in the original paper of SPPO (Wu et al., 2024).
5.3 Response Diversity
We demonstrate an additional advantage introduced by regularization, which is the diversity of the response. We first provide a motivating example with synthetic data in Appendix C.2, which shows that the unregularized self-play may converge to a collapsed response when multiple equally good responses exist. On the contrary, RSPO with maximum entropy regularization can recover all high-reward responses.
@@ 5.3 Response Diversity
The trend of Self-BLEU scores presented in Table 2 (Right) show that applying RSPO with Reverse KL increases response diversity the most, as well as the LCWRs of AlpacaEval 2.0. The application of Forward KL results in slightly less generation diversity than vanilla SPPO, but it still achieves better win rates. The win rates are the highest when Forward KL and Reverse KL are linearly combined for regularization, while the Self-BLEU scores imply that the response diversity is lower than when only Reverse KL is applied. These results highlight that applying regularization in self-play methods can improve test performance and the diversity of the generations simultaneously.
@@ 6 Conclusion
In this paper, we study the regularization in self-play by proposing a framework, namely Regularized Self-Play Policy Optimization (RSPO). Based on RSPO, we can apply different regularization functions for policy updates by adding the regularization term to the loss functions, which is still guaranteed to converge to the NE of the regularized preference optimization game. In the empirical assessments, we achieve significant improvement over the base model and unregularized self-play method, SPPO. We also empirically demonstrate that regularization promotes response diversity. These findings underscore the critical role of regularization as a fundamental component in optimizing self-play alignment.
@@ Appendix A Proofs
In this section, we provide detailed derivations and proofs of propositions.
A.1 Proof of Equivalence between MD and RSPO
@@
Nash-MD and Online Mirror Descent [Munos et al., 2023, Section 6] can be seen as particular instances of our general Regularized Self-Play Policy Optimization (RSPO) (Equation 7).
@@
In this section, we first provide derivations of Nash-MD and Online Mirror Descent to $\mathcal{L}_{\text{RSPO}}$ without external regularization.
Nash-MD practical loss satisfies that
@@
Equation 17 is the definition of practical Nash-MD loss [Munos et al., 2023, Section 7]. Equation 18 holds by adding an subtracting the same element $\log\pi_{t}(y)$. Equation 19 holds due to $\mathbb{E}_{y^{\prime}\sim\pi_{t}^{\mu}}[\mathbb{P}(y\succ y^{\prime})]=\mathbb{P}(y\succ\pi_{t}^{\mu})$. Equation 20 holds since in each iteration before updating, while computing the loss, $y\sim\pi_{\theta}$ is equivalent to $y\sim\pi_{t}$. The learning rate $\eta$ is originally omitted in the paper [Munos et al., 2023]. Here Nash-MD is generalized by $\mathcal{L}_{\text{RSPO}}$ with $\eta=\tfrac{1}{\tau}$ and $R=0$.
@@
A.2 Proof of the Existence of Nash Equilibrium
Proposition A.2.
Nash Equilibrium in the regularized game in Equation 5 exists, and is unique.
Proof.
@@ A.2 Proof of the Existence of Nash Equilibrium
Assumption A.3 (Relative Convexity w.r.t. $\psi$).
We assume the regularization function $R$ of policy $\pi$ is a $1$-strongly convex relative to some function $\psi$. In other words, $\forall\pi,\pi^{\prime}\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$,
@@
We prove that RSPO in Equation 16 is equivalent to GMMD up to multiplying a constant to the gradient, leading to a regularized Nash Equilibrium.
@@
$\nabla_{\theta}D_{\text{KL}}(\pi||\mu)$ $=\nabla_{\theta}\sum_{y}\pi_{\theta}(y)\cdot(\log\pi_{\theta}(y)-\log\mu(y))=\sum_{y}\nabla_{\theta}\pi_{\theta}(y)\cdot(\log\pi_{{\theta}}(y)-\log\mu(y))+\sum_{y}\nabla_{\theta}\pi_{\theta}(y)$
@@
The first equation holds the following directly from the definition of the KL divergence. The second equation holds due to applying the product rule of differentiation. The third equation holds due to simple algebra, and the second term will then vanish because of the sum of the probabilities. The fourth equation holds because of simple algebra. ∎
A.6 Proof of Proposition B.2
@@
$\nabla_{\theta}D_{\text{KL}}(\mu||\pi)$ $=\nabla_{\theta}\sum_{y}\mu(y)\cdot(\log\mu(y)-\log\pi_{\theta}(y))=-\sum_{y}\mu(y)\nabla_{\theta}\log\pi_{\theta}(y)=-\sum_{y}\pi_{\theta}(y)\frac{\mu(y)}{\pi_{\theta}(y)}\nabla_{\theta}\log\pi_{\theta}(y)$
@@
Appendix B Additional Details

### v2

@@ 4.3 Response Diversity and Other Aspects
Table 3: Response diversity of SPPO and RSPO evaluated with Self-BLEU score. The regularization temperatures of RSPO are the same as in Figure 2. RSPO methods involving reverse KL regularization have both better generation quality and diversity.
We demonstrate additional advantages introduced by regularization. We first investigate the diversity of the response cause by regularization. We first provide a motivating example with synthetic data in Appendix D.2, which shows that the unregularized self-play may converge to a collapsed response when multiple equally good responses exist. On the contrary, RSPO with maximum entropy regularization has multi-modal distribution for generation.
@@
The trend of Self-BLEU scores presented in Table 3 (Right) show that applying RSPO with Reverse KL increases response diversity the most, as well as the LCWRs of AlpacaEval 2.0. Although reverse KL regularization is typically associated with reduced diversity, it can, counterintuitively, enhance diversity when the high-probability region of the reference policy $\mu$ contains multiple modes—a scenario commonly arising when $\mu$ is pretrained on a diverse dataset. In such cases, the sampling-regularized optimization process with reverse KL can also induce additional modes in the learned policy distribution, thereby promoting greater diversity in responses. In contrast, IS-Forward KL yields slightly lower diversity, as its importance sampling–based implementation necessitates hard clipping for numerical stability. Compared to reverse KL, the $\chi^{2}$ divergence functions as a stronger regularizer (Huang et al., 2024), promoting diversity, albeit at a slower rate.
Finally, we assess additional aspects of response quality on the Ultrafeedback validation set using ArmoRM Wang et al. (2024b) (Table 4). Reverse KL regularization improves both truthfulness and helpfulness. Notably, while forward and reverse KL regularization individually tend to diminish instruction-following performance, their combination yields improvements across nearly all evaluated aspects, attaining the highest overall score—with particularly strong gains in Instruction Following and Truthfulness.
Methods Overall Score Instruction Following Truthfulness Honesty Helpfulness
Snorkel 0.706 0.781 0.796 0.821 0.760
SPPO 0.716 0.798 0.812 0.836 0.771
RSPO ($\chi^{2}$, $\lambda=0.1$) 0.713 0.793 0.805 0.827 0.769
RSPO (Rev. $\lambda=0.5$) 0.718 0.798 0.805 0.831 0.773
RSPO (Rev. $\lambda=1$) 0.715 0.798 0.807 0.826 0.769
RSPO (For. $\lambda=0.1$) 0.711 0.795 0.809 0.824 0.760
RSPO (For. $\lambda=0.5$) 0.713 0.793 0.815 0.826 0.749
RSPO (For.+Rev.) 0.719 0.805 0.816 0.833 0.768
Table 4: ArmoRM Evaluation. Evaluation of diverse response quality aspects using ArmoRM on Ultrafeedback validation set. The combined application of forward and reverse KL regularization leads to superior performance compared to either form of regularization applied independently.
5 Related Work
Offline RLHF with general divergence for regularization. The use of general divergence-based regularization has been explored in the context of offline alignment. $f$-DPO (Wang et al., 2023) extends Direct Preference Optimization (Rafailov et al., 2024) from reverse KL regularization to a broader class of $f$-divergences, but primarily demonstrates benefits in generation diversity. The specific effects of individual divergences—and their performance on widely-used benchmarks such as AlpacaEval—remain unexamined. $\chi$PO (Huang et al., 2024) emphasizes the theoretical importance of $\chi^{2}$ divergence for uncertainty quantification. However, the role of regularization in online iterative preference optimization, particularly its empirical impact on standard benchmarks, has yet to be studied.
Contrastive Self-Play Alignment We emphasize the distinction between our self-play approach and contrastive self-play methods including Direct Nash Optimization (DNO) (Rosset et al., 2024) and Iterative Nash Policy Optimization (INPO) (Zhang et al., 2024). These methods conduct policy optimization with a loss objective necessary but not sufficient for Mirror Descent (MD) update (Beck and Teboulle, 2003). This objective is constructed via winner-loser response comparisons similar to Direct Preference Optimization (DPO) and Identity Preference Optimization (IPO) (Azar et al., 2024). Optimizing such contrastive loss can lead to only an increase in the relative likelihood gap without necessarily enhancing the absolute probability of the preferred response (Pal et al., 2024). In contrast, our method directly approximates the MD update by converting it to an equivalent reinforcement learning problem, thereby circumventing the limitations of contrastive approaches.
@@ 6 Conclusion
In this paper, we study the regularization in self-play by proposing a framework, namely Regularized Self-Play Policy Optimization (RSPO). Based on RSPO, we can apply different regularization functions for policy updates by adding the regularization term to the loss functions, which is still guaranteed to converge to the NE of the regularized preference optimization game. In the empirical assessments, we achieve significant improvement over the base model and unregularized self-play method, SPPO. We also empirically demonstrate that regularization promotes response diversity, as well as instruction following, truthfulness, and helpfulness. These findings underscore the critical role of regularization as a fundamental component in optimizing self-play alignment.
@@ Appendix A Proofs
In this section, we provide detailed assumptions, derivations and proofs of propositions.
Assumption A.1 (Relative Convexity of $R$ w.r.t. entropy function).
We assume the regularization function $R$ of policy $\pi$ is a $1$-strongly convex relative to entropy function. In other words, $\forall\pi,\pi^{\prime}\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$, and $\psi(\pi)=\langle\pi,\log\pi\rangle$, we have
Assumption A.1 constrains the class of regularization terms $R$ under which theoretical convergence guarantees can be established. Nonetheless, a broad family of divergences still satisfies this assumption, allowing RSPO to retain convergence properties in a wide range of settings. Among the divergences used in our experiments—including linear combinations—only the forward KL divergence violates this assumption. Interestingly, however, forward KL regularization is empirically observed to reduce response length. To leverage this desirable property while preserving theoretical validity, we propose a linear combination of forward and reverse KL divergences, enabling effective length-controlled generation without sacrificing convergence guarantees, and obtains the best generation quality empirically.
A.1 Proof of the Existence of regularized Nash Equilibrium
@@
Nash Equilibrium in the regularized game in Equation 4 exists, and it is unique.
@@
A.2 Proof of Equivalence between MD and RSPO
Proposition A.3.
Nash-MD and Online Mirror Descent [Munos et al., 2023, Section 6] can be seen as instances of Regularized Self-Play Policy Optimization (RSPO) (Equation 6).
Proof.
In this section, we first provide derivations of how Nash-MD is equivalent to RSPO:
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}=\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu}),B=\frac{1}{2},R=D_{\text{KL}}(\pi_{\theta}\|\mu)\big{)}$ (17)
On one hand, Nash-MD practical loss [Munos et al., 2023, Section 7] is defined as
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)$ (18)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{)}\Big{]}$ (19)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}+2\tau\log\frac{\pi_{t}(y)}{\mu(y)}\Big{)}\Big{]}$ (20)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}+\tau\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{)}\Big{]}$ (21)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\Big{)}\Big{]}+\tau\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\log\frac{\pi_{\theta}(y)}{\mu(y)}\Big{]}$ (22)
$=$ $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y)\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\Big{)}\Big{]}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu)$ (23)
$=$ $2\tau^{2}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\nabla_{\theta}\Big{(}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\frac{1}{2\tau}\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}\Big{)}\Big{)}^{2}\Big{]}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu)$ (24)
$=$ $2\tau^{2}\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{t},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\Big{[}\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}-\frac{1}{2\tau}\Big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}\Big{)}\Big{]}^{2}+\tau\nabla_{\theta}D_{\text{KL}}(\pi_{\theta}\|\mu).$ (25)
Equation 19 is the definition of Nash-MD policy gradient. Equation 20 holds because the additional term satisfies that $\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\big{[}\nabla_{\theta}\log\pi_{\theta}(y)\big{(}\log\frac{\pi_{t}(y)}{\mu(y)}\big{)}\big{]}=\nabla_{\theta}\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\theta},\\ y^{\prime}\sim\pi^{\mu}_{t}\end{subarray}}\big{[}\log\frac{\pi_{t}(y)}{\mu(y)}\big{]}=\nabla_{\theta}\log\frac{\pi_{t}(y)}{\mu(y)}=0$. Equation 23 holds due to the definition of reverse KL divergence. Equation 24 is derived by computing the integral of $\log\pi_{\theta}(y)\big{(}\mathbb{P}(y\succ\pi_{t}^{\mu})-\frac{1}{2}-2\tau\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}\big{)}$.
On the other hand, we show that Nash-MD and OMD can also be generalized by RSPO without external regularization, such that we can add additional regularization to existing regularized self-play methods. Nash-MD practical loss [Munos et al., 2023, Section 7] is defined as
@@ A.2 Proof of Equivalence between MD and RSPO
Equation 26 is the definition of practical Nash-MD loss [Munos et al., 2023, Section 7]. Equation 27 holds by adding an subtracting the same element $\log\pi_{t}(y)$. Equation 28 holds due to $\mathbb{E}_{y^{\prime}\sim\pi_{t}^{\mu}}[\mathbb{P}(y\succ y^{\prime})]=\mathbb{P}(y\succ\pi_{t}^{\mu})$. The learning rate $\eta$ is originally omitted in the paper [Munos et al., 2023]. Here Nash-MD is generalized by $\mathcal{L}_{\text{RSPO}}$ with $\eta=\tfrac{1}{\tau}$ and $R=0$.
@@
We prove that RSPO in Equation 8 is equivalent to GMMD up to multiplying a constant to the gradient, leading to a regularized Nash Equilibrium. We follow SPPO to replace the samples $y\sim\pi_{\theta}$ with $y\sim\pi_{t}$ directly since they are equivalent while computing the loss before updating, and rewrite the loss equivalent to GMMD:
$\nabla_{\theta}\mathcal{L}_{\text{GMMD}}(\theta)=\nabla_{\theta}\Bigg{(}\frac{1}{2}{\mathbb{E}_{{y\sim\pi_{t}}}\Big{[}-\eta G(y;\pi_{t})+\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}{+\eta B}\Big{]}^{2}}+\tau R(\pi_{\theta},\mu)\Bigg{)}=\frac{1}{2}\nabla_{\theta}\mathcal{L}_{\text{RSPO}}(\theta).$ (41)
Therefore, according to Equation 41, RSPO is the RL implementation of GMMD, since gradients of losses are equivalent up to multiplying a constant. Then we can derive the convergence gaurantee of RSPO.
@@
$\nabla_{\theta}D_{\text{KL}}(\pi||\mu)$ $=\nabla_{\theta}\sum_{y}\pi_{\theta}(y)\cdot(\log\pi_{\theta}(y)-\log\mu(y))$ (47)
$=\sum_{y}\nabla_{\theta}\pi_{\theta}(y)\cdot(\log\pi_{{\theta}}(y)-\log\mu(y))+\sum_{y}\nabla_{\theta}\pi_{\theta}(y)$
@@
The first equation holds because of the definition of KL divergence. The second equation holds due to applying the product rule of differentiation. The third equation holds due to simple algebra, and the second term will then vanish because of the sum of the probabilities. The fourth equation holds because of simple algebra. ∎
A.6 Proof of Proposition C.2
@@
$\nabla_{\theta}D_{\text{KL}}(\mu||\pi)$ $=\nabla_{\theta}\sum_{y}\mu(y)\cdot(\log\mu(y)-\log\pi_{\theta}(y))$ (49)
$=-\sum_{y}\mu(y)\nabla_{\theta}\log\pi_{\theta}(y)=-\sum_{y}\pi_{\theta}(y)\frac{\mu(y)}{\pi_{\theta}(y)}\nabla_{\theta}\log\pi_{\theta}(y)$
@@
Appendix B Additional Related Work
Preference Optimization.
Large Language Models (LLMs) recently have obtained remarkable capabilities to accomplish a range of tasks [Jiang et al., 2023b, Dubey et al., 2024, DeepSeek-AI et al., 2025], generating more desirable and helpful content following the user’s intention. One of the most important methods to align LLMs with human intentions is Reinforcement Learning from Human Feedback (RLHF), maximizing a preference-based reward penalized by a reverse KL regularization term of the LLM policy and a reference model [Christiano et al., 2017, Ouyang et al., 2022, Rafailov et al., 2024, Azar et al., 2024, Xiong et al., 2024]. Since the reference model usually provides safer guidance for policy optimization [Munos et al., 2023], this regularization is crucial in RLHF to prevent over-optimization, which has been extensively studied and extended beyond KL divergence [Wang et al., 2023, Go et al., 2023, Huang et al., 2024]. In this work, we instead study the regularization problems in self-play alignment.
RLHF with General Preference Optimization (Self-Play Alignment). Azar et al. [2024] introduced the first approach for optimizing LLM policy via general preference models. Nash-MD [Munos et al., 2023] pioneered the application of self-play to general preference optimization by framing it as a two-player game. Subsequent methods have either focused on learning the NE of the original unregularized game (e.g. [Swamy et al., 2024, Wu et al., 2024, Rosset et al., 2024, Wang et al., 2024a]) or the NE of a reverse-KL-regularized preference optimization game (e.g. [Munos et al., 2023, Calandriello et al., 2024, Zhang et al., 2024]). In contrast, our work explores a broader class of divergence-based regularization techniques for self-play alignment.
Notably, our RSPO can generalize existing self-play methods. Unregularized self-play methods following the preference-based MWU can all be generalized by $\mathcal{L}_{\text{RSPO}}$ without external regularization, and thus can be regularized by simply adding regularization term to the loss functions. Based on the same exponential update rule as in SPPO, SPO [Swamy et al., 2024] is equivalent to updating policy with the loss in Equation 9. Magnetic Policy Optimization [Wang et al., 2024a], despite incorporating regularization in the policy update, periodically updates $\mu=\pi_{t}$. Consequently, it inherently follows MWU while incorporating multiple policy updates within each iteration, following
Online iterative RLHF. Iterative alignment method incorporates a reliable reward or preference model—including self-play—functions as a self-improving framework by iteratively generating new data using models and optimizing policies based on this data [Schulman et al., 2017, Ouyang et al., 2022, Bai et al., 2022, Touvron et al., 2023, Dong et al., 2024]. Moreover, extending powerful offline methods such as DPO to iterative frameworks has led to significant performance gains [Xu et al., 2023, Liu et al., 2023, Tran et al., 2023, Dong et al., 2024, Calandriello et al., 2024, Pang et al., 2024, Xiong et al., 2024, Guo et al., 2024, Tajwar et al., 2024, Cen et al., 2024, Xie et al., 2024]. In contrast, our work investigates general preference optimization through self-play from a game-theoretic perspective, shifting the objective from conventional RL optimization to the computation of NE.
Appendix C Additional Details